Test marks (%) obtained in a test are normally distributed with a mean of 55% and variance of 400(% squared). What is the probability that a randomly selected student scored above 60%?


1
Expert's answer
2021-05-28T08:55:02-0400

"\mu =55 \\

\sigma = \sqrt{400} = 20 \\

P(X>60) = 1 -P(X<60) \\

=1 -P(Z< \frac{60-55}{20}) \\

= 1 -P(Z<0.25) \\

= 1 -0.5987 \\

= 0.4013"